OpenStudy (anonymous):

solve the equation by completing the square 2x^2 = 3x + 1

OpenStudy (anonymous):

x=14(3+17√),1/4(3−17√) x≈1.7808,−0.2808

OpenStudy (wolf1728):

Unlike SailorGirl's solution which is correct (I'm guessing it was found by the Quadratic Formula), here is how it is done by completing the square: Divide equation by X²coefficient x^2 -1.5x = .5 Move non-X term to the right (it is already there) Take coefficient of x (-1.5) divide by 2 then square it. -.75² = 0.5625 Add 0.5625 to both sides of the equation x² -1.5x +0.5625 = 1.0625 Take the square root of both sides of the equation (x-.75)(x-.75) = 1.0307764064 (x-.75)(x-.75) -1.0307764064 =0 x = .75 -1.0307764064 x = 1.7807764064 Remembering to use the other square root: x = .75 -1.0307764064 x = -0.2807764064 Yes, it's just that simple - LOL